OpenStudy (anonymous):

Pre-Calculus Help. √6x-2 = √4x+4 Not sure how to approach this one.

OpenStudy (anonymous):

First thing you can do is square both sides.

OpenStudy (anonymous):

It's clear to me, you would square both sides and get: \[ |6x-2|=|4x+4| \]

OpenStudy (anonymous):

@Channi Are you there?

OpenStudy (zarkon):

\[\sqrt{x^2}=|x|\] but \((\sqrt{x})^2=x\) so for this problem you need only solve \(6x-2=4x+4\)

OpenStudy (anonymous):

Oh! I think I can figure it out. I didn't really know where to start. Thanks guys!